There are 104 baseball cards total.
Tom has 8 more cards than Mike.
How many cards do Mike and Tom each have?

Answer by edjones(8007)

m=number of Mike's cards.
t=m+8=Tom's cards.
m+m+8=104
2m=96
m=48
t=56
.
Ed